# Chiral Heisenberg Gross-Neveu-Yukawa criticality: honeycomb vs. SLAC fermions

### v1

@@ Abstract
We perform large scale quantum Monte Carlo simulations of the Hubbard model at half filling with a single Dirac cone close to the critical point, which separates a Dirac semi-metal from an antiferromagnetically ordered phase where SU(2) spin rotational symmetry is spontaneously broken. We discuss the implementation of a single Dirac cone in the SLAC formulation for eight Dirac components and the influence of dynamically induced long-range super-exchange interactions. The finite size behavior of dimensionless ratios and the finite size scaling properties of the Hubbard model with a single Dirac cone are shown to be superior compared to the honeycomb lattice. We extract the critical exponent believed to belong to the chiral Heisenberg Gross-Neveu-Yukawa universality class: The critical exponent ${\nu=1.02(3)}$ coincides for the two lattice types once honeycomb lattices of linear dimension ${L\geq 15}$ are considered. In contrast to the SLAC formulation, where the anomalous dimensions are estimated to be ${\eta_{\phi}=0.73(1)}$ and ${\eta_{\psi}=0.09(1)}$, they remain less stable on honeycomb lattices, but tend towards the estimates from the SLAC formulation.
@@ I Introduction
The family of Gross-Neveu-Yukawa (GNY) universality classes capture the complex interplay of critical Dirac fermions coupled to a scalar bosonic fields. It represents the simplest universality of quantum critical points that lack classical analogues and realize the most fundamental scenario of relativistic fermions: the spontaneous acquisition of mass triggered by interactions [1, 2]. Here we focus specifically on the chiral Heisenberg [O(3) or SU(2)] transition [3] with ${N=8}$ Dirac components in $(2\!+\!1)$ dimensions found, e.g., in semi-metal to insulator transitions, which remains hard to control in analytical approaches [4, 5, 6, 7, 8, 9] as well as quantum Monte Carlo (QMC) simulations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Where the critical properties of the chiral Ising GNY transition have been arguably consistently determined to be at archipelago provided by conformal bootstrap [24], competing interaction channels and generically large corrections in finite size scaling prevent the same for the chiral Heisenberg fixed point [8].
The fairly disparate state of estimates for the ${N=8}$ critical exponents which allow us to classify the universality of the chiral Heisenberg transition is summarized in Fig. 1 along with the estimates from this investigation. In particular the results from QMC simulations, which rightfully pride themselves to be numerically exact, are far from a consensus. This illustrates the ambiguities in the finite size scaling analysis and the limitations by practically accessible spacetime volumes.
Figure 1: Comparison of estimates for the correlation-length exponent $\nu$, the boson anomalous dimension $\eta_{\phi}$ and the fermion anomalous dimension $\eta_{\psi}$ from analytical approaches [4, 5, 6, 7, 8, 9] and Monte Carlo simulations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] for the ${N=8}$ chiral-Heisenberg universality class in chronological order. Gray vertical bars indicate the estimates and their errorbars from the SLAC-Hubbard model discussed in this manuscript.
@@
In the following Sec. II we describe the implementation and simulation details of congruent Dirac cones with a particular number $N$ of zero modes on the lattice and address the reliability and potential issues such as locality, renormalizability and dynamically induced non-local interactions in Sec. III. A finite size scaling analysis and estimates for the critical exponents are presented in Sec. IV. In the following we indicate results obtained with the SLAC formulation with a [Uncaptioned image] -icon, where the inscribed number denotes the number of massless modes $N$. Results obtained on the honeycomb lattice (${N=8}$) are identified with a [Uncaptioned image] -icon.
@@ II.1 Formulation for N-pole SLAC fermions
where ${\sigma=1,\ldots,N_{f}}$ denotes the fermion flavor, ${\bm{\sigma}=(\bm{\sigma}_{x},\bm{\sigma}_{y})}$ is a vector of Pauli matrices and the spinor ${\bm{\psi}^{\dagger}_{\mathbf{k}\sigma}=(a^{\dagger}_{\mathbf{k}\sigma},b^{\dagger}_{\mathbf{k}\sigma})}$ consists of the creation operators at individual orbitals with momentum $\mathbf{k}$. On a square lattice of linear system size ${L=2N_{k}+1}$, where ${-N_{k},-N_{k}+1,\ldots,N_{k}\in\mathbb{N}}$, with two orbitals per unit cell, the primitive vectors in the $x$- and $y$-directions and unit lattice constant, the Fourier transform ${a_{\mathbf{r}\sigma}=\sum_{\mathbf{k}}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}\sigma}/L}$ then results in the bipartite real-space hopping Hamiltonian
@@
[…]
@@
SLAC fermions on lattices with even linear dimension ${L=2N_{k}}$, have the amplitudes  [^1] Contrary to Ref. 51, lattices with even linear dimension do not induce Gibbs ringing, but also exhibit the perfect linear dispersion once the appropriate Fourier transformation for has been performed as shown in App. A.
@@
As required by time reversal symmetry, adiabatic motion of quasiparticles on the Fermi surface as their momentum goes around a loop $\mathcal{C}$ encompassing a band crossing point accumulates integer Berry flux, i.e., ${-\mathrm{i}\oint_{\mathcal{C}}\mathrm{d}\mathbf{k}\cdot\langle\psi_{\mathbf{k}}|\nabla_{\mathbf{k}}|\psi_{\mathbf{k}}\rangle=n\pi}$, where $|\psi_{\mathbf{k}}\rangle$ is the Bloch wave function [30]. The Dirac cone carries a Berry flux of $\pi$ as illustrated in the pseudo spin map ${\mathbf{k}\to\bm{\tau}_{\mathbf{k}}/|\bm{\tau}_{\mathbf{k}}|}$ of the eigenstates $|\psi_{\mathbf{k}}\rangle$ from momentum space to $S^{2}$ in Fig. 2 (with degeneracy points removed) [31]. The correct Berry flux is essential for flux depending ordered states, such as quantum Hall order, but not necessarily for the existence of a phase transitions upon introduction of interactions. The dispersion, as well as the Berry flux are discontinuous at the Brillouin zone (BZ) boundary, which allows to place a single Dirac cone in the first BZ, but constitutes a potential source of singularities, or zeros in the interacting spectrum. Let us note, that any odd $L$ systems does not include the removable discontinuity at the boundary. As shown in Fig. 4, in the interacting system however, the dispersion across the periodic boundaries is renormalized and becomes smooth.
[…]
@@
with ${\bm{\Gamma}=(\bm{\sigma}_{0}\otimes\bm{\sigma}_{x},\bm{\sigma}_{0}\otimes\bm{\sigma}_{y})}$ and the 4-spinor ${\bm{\psi}^{\dagger}_{\sigma}}$ then is the ${N=8}$ generalization of Eq. (1). This formulation results two congruent Dirac cones with $N$ zero modes in total, Berry flux $\pi$ each and can be augmented by the Hubbard interaction
@@ II.2 Method
We employ a projector axillary-field determinantal QMC scheme by which we obtain zero temperature momentum and spin resolved single-particle Green’s function in imaginary time for finite lattices with ${NL^{2}/2}$ sites for SLAC fermions and $2L^{2}$ sites for the honeycomb lattice with periodic boundary conditions. We stochastically project out the ground state from the half-filled slater determinant of the non-interacting system using a projection length ${2\Theta=70}$ and imaginary-time step ${\Delta\tau=0.1}$. An imaginary-time evolution length ${\tau=30}$ has been used to obtain the Green’s function. The Hamiltonian of Eq. (6) is free from a negative sign problem in QMC simulations [36] and the simulation with SLAC fermions allows for local fast updates. Details of the algorithm can be found in Ref. [37].
@@ III Locality & dynamically induced interactions
In this manuscript we simulate the Hamiltonian lattice model in Eq. (6), rather than the GNY lattice field theory itself. We have to be mindful of the apparent long range structure of the Dirac operator in real space, which stands in stark contrast to the inherent locality of physical quantum field theories, but allows to avoid the Nielsen-Ninomiya theorem [38, 39, 40] on finite size systems and to place a single Dirac cone onto the lattice  [^2] Note that in odd spacetime dimensions considered in this work chiral symmetry is replaced by parity symmetry [Winkler15].. The real-space hopping however is not equivalent to a genuine long range coupling, as the hopping only runs along the major axes. This couples a given site only to a finite fraction of the total volume. While this does not exclude possible spurious long-range effects, it suggests their potential influence scales intensively and decreases with increasing system size and dimensionality.
SLAC fermions have been successfully employed in simulations without gauge fields, such as supersymmetric Yukawa, Wess-Zumino and Thirring models [42, 43, 44, 45, 46, 47, 48]. The exactly known critical exponents of the supersymmetric ${N=2}$ chiral-XY transition were accurately reproduced by spinless SLAC fermions [5, 33]. In addition we point to the good agreement between the chiral Ising critical properties from simulations with a single Dirac cone [28] and the recent results from conformal bootstrap [24]. While the chiral Heisenberg transition could be exceptional, the estimates from the other chiral transitions using SLAC fermions are encouraging and do not exhibit pathological signatures.
@@ III.1 Mermin-Wagner criterion & finite temperature
[…]
@@
For ${\alpha<4}$, the energy of the droplet thus grows as a power law, rendering them unlikely and order prevails. The consequence is that the emergence of order which breaks a continuous symmetry is perfectly cromulent in ${2d}$ at finite temperature as long as the spin exchange coupling decays slower than ${J\sim r^{-4}}$. Shorter ranged couplings, where ${\alpha>4}$, result in a bounded droplet energy leading to disorder at finite temperature.
[…]
@@
where ${\alpha<3}$ suffices to allow for the energy to grow, or to allow for the spontaneous breaking of continuous symmetry in ${2d}$ at finite temperature. Consequently, for ${J\sim 1/r^{2}}$ the HMW theorem does not apply, at least within this leading order approximation. Higher order processes will also include spin exchange contributions off the major-axis, but these will likely decay significantly faster than along the major axis. The consequence is, that long-range order can occur within a SLAC-Hubbard model due to the spontaneous breaking of a continuous symmetry in $d<3$ even at finite temperature without violating the HMW theorem. The finite temperature transition is not to be expected in the chiral Heisenberg universality class due to the small exponent ${\alpha=2}$ and the limited dimensionality.
@@ III.2 Zero temperature
At temperature ${T=0}$ the HMW theorem certainly no longer restricts continuous symmetry breaking, as we enter a dimensionality ${d+z}$, where $z$ denotes the dynamic scaling exponent, relating the scaling between the energy and momentum. In particular, Lorentz- (or more specifically Poincaré-) invariance emerges with ${z=1}$ at the critical point in local as well as in SLAC models [27, 28]. However, the dynamically induced long-range super-exchange, if not forbidden by symmetries, will still be present. Here, we investigate its influence and the properties of the Dirac semi-metal (DSM) at weak coupling, as well as the strong coupling AFM and the critical properties.

### v2

@@ Abstract
We perform large scale quantum Monte Carlo simulations of the Hubbard model at half filling with a single Dirac cone close to the critical point, which separates a Dirac semi-metal from an antiferromagnetically ordered phase where SU(2) spin rotational symmetry is spontaneously broken. We discuss the implementation of a single Dirac cone in the SLAC formulation for eight Dirac components and the influence of dynamically induced long-range super-exchange interactions. The finite size behavior of dimensionless ratios and the finite size scaling properties of the Hubbard model with a single Dirac cone are shown to be superior compared to the honeycomb lattice. We extract the critical exponent believed to belong to the chiral Heisenberg Gross-Neveu-Yukawa universality class: The critical exponent $\nu=1.02(3)$} coincides for the two lattice types once honeycomb lattices of linear dimension L>15{L>15} are considered. In contrast to the SLAC formulation, where the anomalous dimensions are estimated to be $\eta_{\phi}=0.73(1)$} and $\eta_{\psi}=0.09(1)$}, they remain less stable on honeycomb lattices, but tend towards the estimates from the SLAC formulation.
@@ I Introduction
The family of Gross-Neveu-Yukawa (GNY) universality classes capture the complex interplay of critical Dirac fermions coupled to a scalar bosonic fields. It represents the simplest universality of quantum critical points that lack classical analogues and realize the most fundamental scenario of relativistic fermions: the spontaneous acquisition of mass triggered by interactions [1, 2]. Here we focus specifically on the chiral Heisenberg [O(3) or SU(2)] transition [3] with N=8{N=8} Dirac components in (2+1)(2\!+\!1) dimensions found, e.g., in semi-metal to insulator transitions, which remains hard to control in analytical approaches [4, 5, 6, 7, 8, 9] as well as quantum Monte Carlo (QMC) simulations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Where the critical properties of the chiral Ising GNY transition have been arguably consistently determined to be at archipelago provided by conformal bootstrap [24], competing interaction channels and generically large corrections in finite size scaling prevent the same for the chiral Heisenberg fixed point [8].
The fairly disparate state of estimates for the N=8{N=8} critical exponents which allow us to classify the universality of the chiral Heisenberg transition is summarized in Fig. 1 along with the estimates from this investigation. In particular the results from QMC simulations, which rightfully pride themselves to be numerically exact, are far from a consensus. This illustrates the ambiguities in the finite size scaling analysis and the limitations by practically accessible spacetime volumes.
Figure 1: Comparison of estimates for the correlation-length exponent $\nu$, the boson anomalous dimension $\eta_{\phi}$ and the fermion anomalous dimension $\eta_{\psi}$, assuming the dynamic scaling exponent z=1{z=1}, from analytical approaches [4, 5, 6, 7, 8, 9] and Monte Carlo simulations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] for the N=8{N=8} chiral-Heisenberg universality class in chronological order. Gray vertical bars indicate the estimates and their errorbars from the SLAC-Hubbard model discussed in this manuscript.
@@
In the following Sec. II we describe the implementation and simulation details of congruent Dirac cones with a particular number $N$ of zero modes on the lattice and address the reliability and potential issues such as locality, renormalizability and dynamically induced non-local interactions in Sec. III. A finite size scaling analysis and estimates for the critical exponents are presented in Sec. IV. In the following we indicate results obtained with the SLAC formulation with a [Uncaptioned image] -icon, where the inscribed number denotes the number of massless modes $N$. Results obtained on the honeycomb lattice (N=8{N=8}) are identified with a [Uncaptioned image] -icon.
@@ II.1 Formulation for N-pole SLAC fermions
where $\sigma=1,\ldots,N_{f}$} denotes the fermion flavor, $\bm{\sigma}=(\bm{\sigma}_{x},\bm{\sigma}_{y})$} is a vector of Pauli matrices and the spinor $\bm{\psi}^{\dagger}_{\mathbf{k}\sigma}=(a^{\dagger}_{\mathbf{k}\sigma},b^{\dagger}_{\mathbf{k}\sigma})$} consists of the creation operators at individual orbitals with momentum $\mathbf{k}$. On a square lattice of linear system size $L=2N_{k}+1$}, where $N_{k},-N_{k}+1,\ldots,N_{k}\in\mathbb{N}$}, with two orbitals per unit cell, the primitive vectors in the $x$- and $y$-directions and unit lattice constant, the Fourier transform $a_{\mathbf{r}\sigma}=\sum_{\mathbf{k}}\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}\sigma}/L$} then results in the bipartite real-space hopping Hamiltonian
@@
where the amplitudes t(r){t(r)} correspond to the Fourier amplitudes from the transformation of Eq. (1), which decay $\propto 1/r$} in the thermodynamic limit (TDL) $L\to\infty$}. Originally introduced for 3+1dd lattice field theories, SLAC fermions were constructed for odd $L=2N_{k}+1$} finite size systems [25], where
@@
SLAC fermions on lattices with even linear dimension $L=2N_{k}$}, have the amplitudes  [^1] Contrary to Ref. 52, lattices with even linear dimension do not induce Gibbs ringing, but also exhibit the perfect linear dispersion once the appropriate Fourier transformation for has been performed as shown in App. A.
@@
As required by time reversal symmetry, adiabatic motion of quasiparticles on the Fermi surface as their momentum goes around a loop $\mathcal{C}$ encompassing a band crossing point accumulates integer Berry flux, i.e., $\mathrm{i}\oint_{\mathcal{C}}\mathrm{d}\mathbf{k}\cdot\langle\psi_{\mathbf{k}}|\nabla_{\mathbf{k}}|\psi_{\mathbf{k}}\rangle=n\pi$}, where $|\psi_{\mathbf{k}}\rangle$ is the Bloch wave function [30]. The Dirac cone carries a Berry flux of $\pi$ as illustrated in the pseudo spin map $\mathbf{k}\to\bm{\tau}_{\mathbf{k}}/|\bm{\tau}_{\mathbf{k}}|$} of the eigenstates $|\psi_{\mathbf{k}}\rangle$ from momentum space to $S^{2}$ in Fig. 2 (with degeneracy points removed) [31]. The correct Berry flux is essential for flux depending ordered states, such as quantum Hall order, but not necessarily for the existence of a phase transitions upon introduction of interactions. The dispersion, as well as the Berry flux are discontinuous at the Brillouin zone (BZ) boundary, which allows to place a single Dirac cone in the first BZ, but constitutes a potential source of singularities, or zeros in the interacting spectrum. Let us note, that any odd $L$ systems does not include the removable discontinuity at the boundary. As shown in Fig. 4, in the interacting system however, the dispersion across the periodic boundaries is renormalized and becomes smooth.
The number of poles of the Dirac operator $N$ is renormalization group relevant at the chiral transition [32]. In order to make a proper comparison of models with SLAC fermions with local hopping lattices, such as the $\pi$}-flux, or honeycomb lattice, the degeneracy of the fermions must be taken into account. While this does not affect the general properties of simulations with SLAC fermions, the critical properties will be different. The most basic Dirac cone in 2+1dd can be formulated for $N=2$ zero modes, i.e., spinless (spin-polarized) fermions with a single flavor $N_{f}$. While $N_{f}=1$ allows to study the chiral-Ising [28] and chiral-XY [33] quantum phase transitions at least two flavors per orbital are required to realize a system where SU(2) symmetry can be spontaneously broken upon the introduction of, e.g, a local Coulomb repulsion. The $2\times 2$} representation of Eq. (2) can be easily generalized by increasing the number of flavors. For $N=4,N_{f}=2$} this has been realized in [27, 21] to investigate the chiral-Heisenberg transition with a single Dirac cone. For $N_{f}>2$}, in the adjoint representation, which is suitable for sign-problem-free AF-QMC simulations [34, 35, 36, 37], this however increases the symmetry of the fermions to $\mathrm{SU}(N_{f})\times\mathrm{SU}(N_{f})$} and a naive generalization of the Hubbard interaction would tend to break the $\mathrm{SU}(N_{f})$ symmetry, rather than the targeted SU(2) flavor symmetry. In order to study the chiral-Heisenberg transition, e.g., with N=8{N=8} either the interaction terms must be chosen, such that solely the SU(2) subgroup is spontaneously broken. Alternatively, one may set the number of orbitals $N_{\ell}=4$} per site while keeping $N_{f}=2$: The $4\times 4$} Hamiltonian
@@
with $\bm{\Gamma}=(\bm{\sigma}_{0}\otimes\bm{\sigma}_{x},\bm{\sigma}_{0}\otimes\bm{\sigma}_{y})$} and the 4-spinor $\bm{\psi}^{\dagger}_{\sigma}$} then is the N=8{N=8} generalization of Eq. (1). This formulation results two congruent Dirac cones with $N$ zero modes in total, Berry flux $\pi$ each and can be augmented by the Hubbard interaction
@@ II.2 Method
We employ a projector axillary-field determinantal QMC scheme by which we obtain zero temperature momentum and spin resolved single-particle Green’s function in imaginary time for finite lattices with $NL^{2}/2$} sites for SLAC fermions and $2L^{2}$ sites for the honeycomb lattice with periodic boundary conditions. We stochastically project out the ground state from the half-filled slater determinant of the non-interacting system using a projection length $2\Theta=70$} and imaginary-time step $\Delta\tau=0.1$}. An imaginary-time evolution length $\tau=30$} has been used to obtain the Green’s function. The Hamiltonian of Eq. (6) is free from a negative sign problem in QMC simulations [36, 37] and the simulation with SLAC fermions allows for local fast updates. Details of the algorithm can be found in Ref. [38].
@@ III Locality & dynamically induced interactions
In this manuscript we simulate the Hamiltonian lattice model in Eq. (6), rather than the GNY lattice field theory itself. We have to be mindful of the apparent long range structure of the Dirac operator in real space, which stands in stark contrast to the inherent locality of physical quantum field theories, but allows to avoid the Nielsen-Ninomiya theorem [39, 40, 41] on finite size systems and to place a single Dirac cone onto the lattice  [^2] Note that in odd spacetime dimensions considered in this work chiral symmetry is replaced by parity symmetry [85].. The real-space hopping however is not equivalent to a genuine long range coupling, as the hopping only runs along the major axes. This couples a given site only to a finite fraction of the total volume. While this does not exclude possible spurious long-range effects, it suggests their potential influence scales intensively and decreases with increasing system size and dimensionality.
SLAC fermions have been successfully employed in simulations without gauge fields, such as supersymmetric Yukawa, Wess-Zumino and Thirring models [43, 44, 45, 46, 47, 48, 49]. The exactly known critical exponents of the supersymmetric N=2{N=2} chiral-XY transition were accurately reproduced by spinless SLAC fermions [5, 33]. In addition we point to the good agreement between the chiral Ising critical properties from simulations with a single Dirac cone [28] and the recent results from conformal bootstrap [24]. While the chiral Heisenberg transition could be exceptional, the estimates from the other chiral transitions using SLAC fermions are encouraging and do not exhibit pathological signatures.
@@ III.1 Mermin-Wagner criterion & finite temperature
The Hamiltonian lattice model in Eq. (6) and SLAC fermions are by construction effective low energy models akin to an effective field theory. As such their application is confined by energy scales and dimensionality [25]. Indiscriminate application of the SLAC Hamiltonian in a regime, dimensionality, or limit it is not intended for, can lead to physical scenarios which are not compatible with the original local Hamiltonian the effective Hamiltonian has been modelled on. This has been demonstrated in [50, 51, 52] where the effects of long range hopping result in the spontaneous breaking of continuous symmetries in one dimension. In 2dd at finite temperature the authors of [53] report an interaction induced finite temperature phase transition in a SLAC-Hubbard model, which is absent in the purely short-range $\pi$-flux Hubbard model on the square lattice. While at odds with the physics of the short-range model, which forbids the spontaneous breaking of a continuous symmetry, the model with SLAC fermions does not violate the Hohenberg-Mermin-Wagner (HMW) theorem [54, 55], since it does not apply.
Despite claims of a break down, or violation of the HMW theorem in models with fully connected long range interactions [53, 56], for spin-couplings $J\sim r^{-\alpha}$} with $\alpha<4$} the conditions of the theorem are not satisfied: The theorem considers a magnetically ordered 2dd spin system at finite temperature and the energy necessary to create a magnetically disordered droplet of volume $V\sim L^{2}$ with linear size $L$, which is of the order $\sum_{r\in V}r^{2}J(r)$, [57]. If this energy is bounded, independently of the size $V$, the disorder droplets are probable and destroy the magnetic order. For fully (all-to-all) connected long range interactions of the kind $J\sim r^{-\alpha}$}, where $\mathbf{r}=(x,y)$}, the droplet energy is at worst
@@
For $\alpha<4$}, the energy of the droplet thus grows as a power law, rendering them unlikely and order prevails. The consequence is that the emergence of order which breaks a continuous symmetry is perfectly cromulent in 2d{2d} at finite temperature as long as the spin exchange coupling decays slower than $J\sim r^{-4}$}. Shorter ranged couplings, where $\alpha>4$}, result in a bounded droplet energy leading to disorder at finite temperature.
The SLAC formalism, can be imagined as an 2dd bilayer system with all-to-2L{2L} connectivity along the major axis, such that the connectivity scales sub-volume. The SLAC hopping matrix elements $t\sim 1/r$}, generate to leading order the effective AFM exchange coupling $J\sim t^{2}/U\sim 1/Ur^{2}$}. In analogy with the example above, at strong coupling, for major-axis connectivity the disorder droplet energy becomes
@@
where $\alpha<3$} suffices to allow for the energy to grow, or to allow for the spontaneous breaking of continuous symmetry in 2d{2d} at finite temperature. Consequently, for $J\sim 1/r^{2}$} the HMW theorem does not apply, at least within this leading order approximation. Higher order processes will also include spin exchange contributions off the major-axis, but these will likely decay significantly faster than along the major axis. The consequence is, that long-range order can occur within a SLAC-Hubbard model due to the spontaneous breaking of a continuous symmetry in $d<3$ even at finite temperature without violating the HMW theorem. The finite temperature transition is not to be expected in the chiral Heisenberg universality class due to the small exponent $\alpha=2$} and the limited dimensionality.
@@ III.2 Zero temperature
At temperature T=0{T=0} the HMW theorem certainly no longer restricts continuous symmetry breaking, as we enter a dimensionality d+z{d+z}, where $z$ denotes the dynamic scaling exponent, relating the scaling between the energy and momentum. In particular, Lorentz- (or more specifically Poincaré-) invariance emerges with z=1{z=1} at the critical point in local and in SLAC models [58, 27, 28], which can be identified from the scaling of the single particle gap $\Delta_{\text{sp}}\sim L^{-z}$} at the critical coupling in the inset of Fig. 3. However, the dynamically induced long-range super-exchange, if not forbidden by symmetries, will still be present. Here, we investigate its influence and the properties of the Dirac semi-metal (DSM) at weak coupling, as well as the strong coupling AFM and the critical properties.